OpenStudy (anonymous):

Let g(theta)= (theta)cos(theta). Find g'(pi/3). I'm confused about which derivative rule I should be using. Quotient? I may not be approaching the problem correctly at all.

OpenStudy (anonymous):

I'll give it a try that way, thank you!

OpenStudy (anonymous):

Product Rule work.

OpenStudy (owlcoffee):

the product rule is: \[(f.g)'(x)=f'(x)g(x)+f(x)g'(x)\]

OpenStudy (anonymous):

product rule\[1\cos (\theta)-\theta*\sin(\theta) =1/2-\sqrt{3}/2\]